OpenStudy (anonymous):

PLEASE HELP Solve the equation. x^2 + 12x + 36 = 25 (1 point) 11, 1 –1, 1 –1, –11 11, –11

OpenStudy (solomonzelman):

What methods do you know? Quadratic formula? Completing the square? Factoring?

OpenStudy (anonymous):

factoring

OpenStudy (solomonzelman):

What about completing the square, b/c the equation is like designed for that!

OpenStudy (solomonzelman):

For factoring, your first step would be, to subtract 25 from both sides.

OpenStudy (anonymous):

so im subtracting 25 from 36

OpenStudy (solomonzelman):

yes!

OpenStudy (solomonzelman):

(and on the right side, 25 from 25)

OpenStudy (anonymous):

ok then my problem comes to x^2+12x=11

OpenStudy (solomonzelman):

No! it is -11, b/c originally when you subtract you get x^2+12x+11=0. See?

OpenStudy (anonymous):

o ok

OpenStudy (solomonzelman):

So your equation is \[x^2+12x+11=0.\]

OpenStudy (solomonzelman):

Can you factor that?

OpenStudy (solomonzelman):

i can give you couple tips on how to factor this, i am sure it will make it much easier!!

OpenStudy (anonymous):

x^+1=0

OpenStudy (anonymous):

@whiterabbit541

OpenStudy (solomonzelman):

Hold on, that is not right! You need to factor out the \[x^2+12x+11=0\]

OpenStudy (anonymous):

Is the answer a

OpenStudy (solomonzelman):

I'll give you a couple tips on this.... 1) All numbers when it is factored are positive. 2) the constants must be 1 and 11 Still need help o how to factor this? Don't 11 and 1 add to 12?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\[(x+11)(x+1)\] what about now?

OpenStudy (solomonzelman):

\[(x+11)(x+1)=0\] therefore \[x+1=0\] or \[X+11=0\] So what is x?